# Asynchronous expressed-private $q$-voter model on networks: self-anticonformity and preference falsification

### v2

@@ IV.3 Results on empirical networks
We use the organizational networks collected by Fire et al. Fire and Puzis (2016). Although we conducted simulations on all networks reported in Fire and Puzis (2016), here we present results only for two of them: the small network S2, consisting of $N=320$ nodes with average degree $k=14.81$ and average clustering coefficient $C=0.49$, and the medium network M2 with $N=3862$ nodes, $k=45.22$, and $C=0.31$. Average node degrees of the two selected networks are the closest to Dunbar’s number describing sizes of social circles (5, 15, 50 and 150 people who are increasingly distant). Furthermore, structure of these networks lacks hubs (nodes of high degree, whose nodes are sparsely connected with each other), that violate pair approximation assumptions.
@@
Results for the small networks (panels a and d in Fig. 5) are noisy due to finite-size fluctuations. Despite this, we surprisingly find that for the S2 network the simulations on the empirical graph agree better with the analytical predictions than those on its Watts–Strogatz counterpart.
For the M2 network (panels b and e in Fig. 5), the results are, first, much smoother. Second, they agree better with the analytical calculations, since this network has a lower clustering coefficient and thus weaker short-range correlations between neighboring nodes.
Finally, in the last column of Fig. 5 we compare the pair-approximation results with simulations on random graphs of size $N=10^{4}$ and average degree $k=15,50,150$. This network class best matches the pair-approximation assumption of weak correlations between neighbors.
@@
Figure 5: Stationary concentration of positive expressed opinions $c_{S}$ as a function of the independence probability $p$ for $\alpha=0.1,q=3$. Monte Carlo results on empirical networks Fire and Puzis (2016) are compared with simulations on Watts–Strogatz networks matched in size, average degree, and clustering coefficient (panels a, b, d, e). Panels (c, f) show simulations on random graphs of size $N=10^{4}$ with average degree $k=15,50,150$. Full markers correspond to ordered initial conditions, while empty markers correspond to random initial conditions. Simulations results are compared with the pair-approximation prediction. Results are shown for both model variants: with (top row) and without (bottom row) self-anticonformity.
@@ V Conclusions
In this paper, we introduced and compared two versions of the $\alpha$-EPO model, extending the two $q$-voter EPO variants proposed in Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025). These two variants differ in that one includes self-anticonformity whereas the other does not. A parallel analysis of the corresponding $\alpha$-EPO versions allows us to investigate the role of self-anticonformity, which has recently attracted significant attention Dvorak et al. (2024); Lipiecki and Sznajd-Weron (2025a). Our modification of the original models is motivated by the fact that people’s private and expressed opinions, beliefs or attitudes do not necessarily change simultaneously or at the same rate. Depending on the context, these states may represent, for example, private versus expressed opinions or intentions versus behavior. In the $\alpha$-EPO models, agents update their private opinion with probability $\alpha$ and their expressed opinion with complementary probability $1-\alpha$.
@@
Once self-anticonformity is excluded, this robustness disappears: the stationary dependencies $c_{S}(p)$ and $c_{\sigma}(p)$ change with $\alpha$. In this variant, our results approach those of the original model Kamińska and Sznajd-Weron (2025) in the limit $\alpha\to 0$, which can be understood intuitively from the update rules: when updates occur predominantly at the expressed level, the expressed layer relaxes faster, and private conformity events are more likely to be effective because they are permitted only when the target’s expressed opinion already matches the unanimous influence group. For influence-group sizes $q\geq 2$, the critical point $p^{*}$ increases with $\alpha$, meaning that for larger $\alpha$ it is easier to reach agreement; a majority of agents choose the same option even for a higher probability of independence $p$. Moreover, as $\alpha$ increases, the hysteresis region associated with a discontinuous phase transition becomes narrower.
For $q=3$, this effect is particularly clear: as $\alpha$ increases, the hysteresis region not only narrows but disappears entirely at a critical value $\alpha^{*}$, indicating a change from a discontinuous to a continuous phase transition. This reduction of hysteresis weakens bistability and path dependence, making collective outcomes less sensitive to initial conditions and less prone to delayed switching when conditions change. Thus, from the perspective of the whole group, in contrast to the individual level, it is advantageous when private attitudes are more volatile than publicly expressed behavior. This scenario is also consistent with psychological findings suggesting that people may change their attitudes more frequently than their behavior Kroesen and Chorus (2018).
We show that the above observations hold not only on the complete graph (all-to-all interactions) but also on sparser artificial and empirical networks. Based on these results, we can draw implications that may be relevant for managers and policy makers. First, it may be beneficial to encourage some people to play the devil’s advocate Gerstel et al. (2025) and openly disagree with the majority. Second, it may be useful to discourage overly frequent changes in publicly expressed behavior. If open disagreement is encouraged and publicly expressed behavior is less volatile, agreement may become easier to reach and delayed responses to dynamically changing situations may be reduced. These implications should be interpreted with caution, as they are derived from a specific class of models. Extending this type of analysis to other EPO models is an important direction for future work to assess the generality of our conclusions and to better understand the role of self-anticonformity.

### v3

@@ IV.3 Results on empirical networks
We use the organizational networks collected by Fire et al. [7]. Although we conducted simulations on all networks reported in [7], here we present results only for two of them: the small network S2, consisting of $N=320$ nodes with average degree $k=14.81$ and average clustering coefficient $C=0.49$, and the medium network M2 with $N=3862$ nodes, $k=45.22$, and $C=0.31$. Average node degrees of the two selected networks are the closest to Dunbar’s number describing sizes of social circles (5, 15, 50 and 150 people who are increasingly distant). Furthermore, structure of these networks lacks hubs (nodes of high degree, whose nodes are sparsely connected with each other), that violate pair approximation assumptions.
@@
Results for the small networks (panels a and d in Fig. 5) are noisy due to finite-size fluctuations. Despite this, we surprisingly find that for the S2 network the simulations on the empirical graph agree better with the analytical predictions than those on its Watts–Strogatz counterpart. For the M2 network (panels b and e in Fig. 5), the results are, first, much smoother. Second, they agree better with the analytical calculations, since this network has a lower clustering coefficient and thus weaker short-range correlations between neighboring nodes. Finally, in the last column of Fig. 5 we compare the pair-approximation results with simulations on random graphs of size $N=10^{4}$ and average degree $k=15,50,150$. This network class best matches the pair-approximation assumption of weak correlations between neighbors.
@@
Figure 5: Stationary concentration of positive expressed opinions $c_{S}$ as a function of the independence probability $p$ for $\alpha=0.1,q=3$. Monte Carlo results on empirical networks [7] are compared with simulations on Watts–Strogatz networks matched in size, average degree, and clustering coefficient (panels a, b, d, e). Panels (c, f) show simulations on random graphs of size $N=10^{4}$ with average degree $k=15,50,150$. Full markers correspond to ordered initial conditions, while empty markers correspond to random initial conditions. Simulations results are compared with the pair-approximation prediction. Results are shown for both model variants: with (top row) and without (bottom row) self-anticonformity.
@@ V Conclusions
In this paper, we introduced and compared two versions of the $\alpha$-EPO model, extending the two $q$-voter EPO variants proposed in [13, 16]. These two variants differ in that one includes self-anticonformity whereas the other does not. A parallel analysis of the corresponding $\alpha$-EPO versions allows us to investigate the role of self-anticonformity, which has recently attracted significant attention [5, 22]. We modified the original models by allowing internal and public states to evolve at different rates, which is motivated by the fact that people’s private and expressed opinions, beliefs and attitudes, or intentions and behaviors do not necessarily change simultaneously or equally often.
@@
Once self-anticonformity is excluded, this robustness disappears: the stationary dependencies $c_{S}(p)$ and $c_{\sigma}(p)$ change with $\alpha$. In this variant, our results approach those of the original model [16] in the limit $\alpha\to 0$, which can be understood intuitively from the update rules: when updates occur predominantly at the expressed level, the expressed layer relaxes faster, and private conformity events are more likely to be effective because they are permitted only when the target’s expressed opinion already matches the unanimous influence group. For influence-group sizes $q\geq 2$, the critical point $p^{*}$ increases with $\alpha$, meaning that for larger $\alpha$ it is easier to reach an agreement; a majority of agents choose the same option even for a higher probability of independence $p$. Moreover, as $\alpha$ increases, the hysteresis region associated with a discontinuous phase transition becomes narrower.
For $q=3$, this effect is particularly clear: as $\alpha$ increases, the hysteresis region not only narrows but disappears entirely at a critical value $\alpha^{*}$, indicating a change from a discontinuous to a continuous phase transition. This reduction of hysteresis weakens bistability and path dependence, making collective outcomes less sensitive to initial conditions and less prone to delayed switching when conditions change. Thus, from the perspective of the whole group, in contrast to the individual level, it is advantageous when private attitudes are more volatile than publicly expressed behavior. This scenario is also consistent with psychological findings suggesting that people may change their attitudes more frequently than their behavior [19].
We show that the above observations hold not only on the complete graph (all-to-all interactions) but also on sparser artificial and empirical networks. Based on these results, we draw implications that may be relevant for managers and policymakers. First, it may be beneficial to encourage some people to play the devil’s advocate [10] and openly disagree with the majority. This result is particularly important in light of recent experiments in behavioral economics showing that standing out from the group can be a strategic response in situations where individuals are evaluated [5]. Second, it may be useful to discourage overly frequent changes in publicly expressed behavior. If open disagreement is encouraged and publicly expressed behavior is less volatile, agreement may become easier to reach and delayed responses to dynamically changing situations may be reduced.
We are aware that the above conclusions should be treated with caution, as they are derived from a specific class of models. An important direction for future work is to extend this type of analysis to other $\alpha$-EPO models—an umbrella term for models in which the interplay between private and expressed opinions is controlled by the parameter $\alpha$, which tunes the relative update rates of the two layers—in order to assess the generality of our conclusions and to better understand the role of anticonformity.